OpenStudy (austinl):

Find the solution of the given initial value problem in explicit form. \(y\prime=(1-2x)y^2~,~y(0)=\frac{-1}{6}\)

OpenStudy (austinl):

I am not even sure where to begin, it isn't in any form that I am familiar with.

OpenStudy (loser66):

\[\frac{dy}{y^2}= (1-2x)dx\]integral both sides

OpenStudy (austinl):

\(\large\int \dfrac{dy}{y^2}=?\) What? How the...... I feel silly now, but...

OpenStudy (loser66):

yup, \[\int y^{-2}dy=?\]

OpenStudy (austinl):

Aha, Now I see. Yep, I feel silly.

OpenStudy (loser66):

ok,

OpenStudy (loser66):

@CarlosGP watch him out, please. I have to go somewhere to eat something. Cannot stay

OpenStudy (austinl):

\(\large{-\frac{1}{y}+C=x-x^2+C}\) This time I actually need to solve for y, yes?

OpenStudy (anonymous):

If you put C on both sides they will cancel each other:. Put just one C and calculate it using y(0)=-1/6

OpenStudy (austinl):

\(\large{y(x)=\frac{1}{x(x-1)}+C}\)

OpenStudy (anonymous):

either way is ok. leaving y(x) alone is not always possible

OpenStudy (austinl):

That can't be right.

OpenStudy (anonymous):

now do x=0 and get the "C" that makes y(0)=-1/6

OpenStudy (austinl):

But... the denominator can't be zero. I am lost now.

OpenStudy (anonymous):

.Because the solution is: \[y(x)=\frac{ 1 }{ x(x-1)+C }\]

OpenStudy (anonymous):

It is what you get from: \[-\frac{ 1 }{ y }=x-x^2+C\]

OpenStudy (austinl):

\(\large{y(x)=\dfrac{1}{x(x-1)-6}}\) Is this correct?

OpenStudy (anonymous):

that is ok. Now check that\[y'=(1-2x)y^2\]

OpenStudy (austinl):

Actually, \(\large{y(x)=\dfrac{1}{x^2-x-6}}\) And that is the solution that my text book gives! :D Thanks very much guys!

OpenStudy (anonymous):

you are welcome